What is the probability of a couple (XX and XY) having three daughters in a row?
1
Expert's answer
2020-11-06T09:56:05-0500

The sex of the baby is basically determined by the father depending on whether he gives a X or Y gene for the fusion of the gonosome. There is hence a 1 in 2 chance (50% or 1/2 probability) of either a son or a daughter on each occasion.

Therefore, by the multiplication principle, the probability of having three daughters in a row is

"\frac{1}{2} \times \frac{1}{2} \times \frac{1}{2} = \frac{1}{8} = 0.125"